OpenStudy (anonymous):

how much greater is the surface area of a cube when the length of each edge is multiplied by 7

OpenStudy (anonymous):

S=6a^2 if each edge is multiplied by 7 S'=6(7a)^2=49*(7a^2)=49*S

OpenStudy (anonymous):

what if each edge was multiplied by 6

OpenStudy (anonymous):

put instead of a 6a 36 times greater